OpenStudy (anonymous):

The population of a village is 5000. It is increasing at the rate of 10% each year. At the end of 3 years, 1155 people migrate from the village. The net increase in population at the end of 3 years is??

OpenStudy (lukebluefive):

Population in - Year 0: 5,000 Year 1: 5,000 x 1.1 Year 2: 5,000 x 1.1^2 Year 3: 5,000 x 1.1^3 - 1,155 Net increase = Year 3 Population - 5,000

OpenStudy (anonymous):

got it thanq

OpenStudy (anonymous):

Two pipes A and B can separately fill a cistern in 24 and 32 minutes respectively. They started to fill a cistern together, but B is turned off after few minutes and A fills the rest of the cistern. If total time taken is 18 minutes, after how many minutes is tap B turned off?

OpenStudy (lukebluefive):

Pipe A fills the cistern as per the following equation: \[A = \frac{x}{24}; 0 ≤ x ≤ 24\] And Pipe B: \[B = \frac{x}{32}; 0 ≤ x ≤ 32\] Where x is the number of minutes. Combined they work based on the following equation: \[A+B = \frac{x}{24} + \frac{x}{32} = \frac{(32x+24x)}{768} = \frac{56x}{768} = \frac{7x}{96}\] So, they can both fill the cistern in just under 14 minutes (96/7). But we need different values for the number of minutes each was on: \[A+B = \frac{x}{24} + \frac{y}{32} = \frac{(32x+24y)}{768}\] If x = 18, plug in a solve for y.